OpenStudy (anonymous):

if f(x)=3(x+5)+4/x what is f(a+2) a. 3(a+2)+(4/a)+2 b.3(a+7)+4/a+2 c.3(f(a)+5)+4/f(a)+2 Please help ;.;

OpenStudy (anonymous):

In your original function f(x) replace all your x's with (a+2).

OpenStudy (anonymous):

For example, If i have\[f(x) = 2(x+2)\] and I am asked what is f(a+2). Since f is a function of x I will replace all my X values with my new a+2 and get:\[f(a+2) = 2((a+2)+2)\] Use that same method in your problem and replace all your X values with your new f(a+2)

OpenStudy (anonymous):

Okay thank you, so would it be c? because I add the 2 and the 5 right and then i get x+7?

OpenStudy (anonymous):

I mean b*****

OpenStudy (anonymous):

Correct! Great job

OpenStudy (anonymous):

Thank you so much!!!

OpenStudy (anonymous):

My pleasure.